# Point-In-Context: Understanding Point Cloud via In-Context Learning

### v1

@@ 4.2 Point-In-Context-Segmenter Model
TABLE II: Comparison results on Human & Object Segmentation In-Context Datasets, including ShapeNetPart [3], Human3D [4], and BEHAVE [5]. Additionally, we conduct one-shot generalization testing experiments on AKB-48 [6], which is not included in the training set. We report the mIoU metric.
@@
[…]
@@
PointNet [38] CVPR’17 77.45 80.12 81.05 83.91 78.63 /
DGCNN [41] TOG’19 76.16 78.23 74.57 76.28 86.18 /
PCT [48] CVM’21 79.46 83.71 81.65 83.16 79.42 /
@@
PointNet [38] CVPR’17 75.77 79.07 77.44 79.33 66.06 41.38
DGCNN [41] TOG’19 75.16 78.58 70.52 74.55 77.48 36.07
PCT [48] CVM’21 79.65 82.00 79.46 80.75 72.41 40.17
ACT [74] ICLR’23 84.43 85.98 82.58 84.69 88.33 34.69
I2P-MAE [28] CVPR’23 74.84 77.68 74.36 78.57 79.00 16.89
@@
PIC-G-Cat / 64.52 70.50 72.79 75.28 83.66 9.97
PIC-G-Sep / 79.22 81.72 74.74 77.93 84.03 19.58
PIC-S-Cat / 84.13 87.31 80.11 83.34 88.03 73.52
PIC-S-Sep / 85.14 87.82 82.82 85.59 88.63 63.08
Inference of PIC-S. In contrast to PIC-G’s fixed segmentation target label point map, the in-context labeling training strategy introduces variability in the label point map for each point cloud. This difference poses unique challenges during evaluation, particularly when the target point cloud for the query is not visible. To address this issue, we leverage the label point bank $B_{j}$ utilized by the prompt $Q_{j}$ as the segmentation standard for the query.
During evaluation, for each point in the predicted point cloud, we compute its distance from each point in the label point set $B_{j}$. By selecting the label point with the minimum distance, we determine the segmentation result for each point. This process enables us to transform the dynamic regressed cloud $\hat{P}_{i}$ into corresponding classification prediction while avoiding the fixed property from classification.
Loss Function. As the Chamfer Distance [104] (CD) loss only measures the similarity between two point clouds, in our segmentation benchmark, we require precise regression of predictions for each point. Therefore, we incorporate Smooth-$\ell_{1}$ [105] loss to assist the model in optimization. At the same time, our training purpose is to minimize the distance between the predicted point cloud and the target point cloud, so the CD loss in PIC-G is also used. The loss function of PIC-S can be formulated as:
5 Experiments
We first provide specific experimental details, including model parameters and training procedures, and information on other related methods (Sec. 5.1). Subsequently, we quantitatively and qualitatively evaluate the proposed PIC-G and PIC-S models on the ShapeNet In-Context Dataset (Sec. 5.2) and Human & Object Segmentation In-Context Dataset (Sec. 5.3), comparing them with other methods. The focus lies in exploring their performance in multitasking and part segmentation scenarios. Finally, we conduct detailed ablation studies to examine the rationality of the model architecture design and selection (Sec. 5.4).
Figure 5: Visualization of predictions from PIC-G-Sep on ShapeNet In-Context Datasets. For part segmentation, we visualize the generated target together with the mapping back, both adding category-specific colors for a better look.
5.1 Experimental Details
Implementation Details. We sample $N=1024$ points of each point cloud and divide it into $N_{C}=64$ point patches, each with $M=32$ neighborhood points. Besides, we set the mask ratio as $0.7$. We randomly select a prompt pair that performs the same task with the query point cloud from the training set. We consider $R_{rec}\in\{3\%,6\%,12\%,25\%,50\%\}$ and $R_{den}\in\{10\%,20\%,30\%,40\%,50\%\}$. We use an AdamW optimizer [106] and cosine learning rate decay, with the initial learning rate as 0.001 and a weight decay of 0.05. For PIC-S, We use a pre-trained checkpoint from PIC-G as model initialization. All models are trained for 300 epochs.
Baseline Methods and Training Details To evaluate the proposed framework, we evaluate the two variants (“-Sep” and “-Cat”) of PIC-G and PIC-S on the datasets mentioned in Sec. 3 respectively. We compare them with the following related methods:
1) Point-BERT [67] is a masked auto-encoder. Like the settings in PIC-Cat, we concatenate $2$ pairs of input-target point cloud tokens as a token sequence and input it to Point-BERT, which takes a token sequence as the input and converts it into discrete point tokens from a pre-trained dVAE [66] vocabulary of size 8192.
2) Task-Specific Models. We select three representative methods: PointNet [38], DGCNN [41], PCT [48], and ACT [74], and train them individually on each task or dataset. For the multitasking benchmark, we also design task-specific heads for each task.
3) Multitask Models. For a fair comparison, we develop multitask models based on common point cloud analysis methods. These models feature a shared backbone network and multiple task-specific heads designed to be capable of multitasking learning.
4) PointMAE [27] is a masked auto-encoder. Unlike Point-BERT, PointMAE directly rebuilds points in each local area.
5) ACT [74], I2P-MAE [28], and ReCon [73] are recent SOTA methods that involve other modalities like image and text knowledge in the pre-training stage and enhance the performance on different tasks after fine-tuning the models. Similar to multitask models, we combine a pre-trained encoder with different task heads for simultaneous training on the four tasks.
@@ 5.2 Results on Point Cloud Multitasking
Figure 6: Visualization of comparison results between PIC-G and multitask models on reconstruction (lines 1-2), denoising (lines 3-4), and registration (lines 5-6), where our models can generate more accurate predictions than other multitask models.
Figure 7: (a) Generalized to out-of-distribution data. We evaluate the registration task on the ModelNet40 [32] dataset, which is not present in the training set. The Chamfer Distance (x1000) is used as the evaluation metric. (b) Generalized to new tasks. Our model can rotate at any angle according to the example pairs and show the completion ability to the broken point cloud, which can be treated as an upgraded task of registration.
We report extensive experimental results of various models on the dataset we proposed in Tab. I. From this, we find that our PIC-G and PIC-S exhibit impressive results and are capable of adapting to different tasks after only one training, achieving SOTA results in all four tasks. Besides, we visualize the in-context 3D inference results of PIC-G-Sep in Fig. 5, where our model can generate the corresponding predictions given the provided prompts among all four tasks, including reconstruction, denoising, registration, and part segmentation.
Comparison to Task-Specific Models. Task-specific models, including PointNet [38], DGCNN [41], PCT [48], and ACT [74] outperform PIC-G and PIC-S in most indicators on all tasks. However, in the part segmentation task, PIC-S achieves better results than PointNet, DGCNN, PCT, and ACT, though they are trained for segmentation. It needs to be clarified that the performance of our model is directly related to the choice of prompts. When the quality of the prompt is better, PIC-G and PIC-S can achieve much better results, and this will be demonstrated in the following section.
@@
Figure 8: Visualization of comparison results between two versions of PIC: PIC-S (extended version) and PIC-G (vanilla version). PIC-S generates more accurate segmentation results than PIC-G, as can be seen from the spots with red arrows ($\searrow$) pointed to them.
@@
Figure 9: Generalization of PIC-S. We use customized prompts to guide the model to perform specified part segmentation. The red boxes indicate the output of the PIC-S-Sep.
@@
Comparison to Multitask Models. As shown in Tab. I, our PIC-G and PIC-S outperform multitask models in multitasking learning, especially in part segmentation, while the multitask models suffer from confusion from multiple types of tasks during training. We also conduct a comparison of visualization results between our PIC-G models and multitask models on three tasks, including reconstruction, denoising, and registration. As shown in Fig. 6, compared with other multitask models, our PIC-G-Sep and PIC-G-Cat output results are more satisfactory.
@@
Generalization. Usually, in-context learning has a certain degree of generalization ability, thus allowing the model to adapt to different tasks quickly. This also applies to our models. First, we test the proposed methods on out-of-distribution point clouds by conducting a registration evaluation on the ModelNet40 [32] dataset. As shown in Fig. 7 (a), our models, PIC-G-Sep and PIC-G-Cat, both exhibit superior performance on open-class tasks compared to supervised learning models specifically trained for the registration task. We compare it with models trained on the single task, such as PointMAE [27], and PCT [48]. These models suffer obvious performance drops when transferred to the new dataset. The greater the difficulty, the greater the decline in performance.
@@
TABLE III: Effectiveness of Joint Sampling Module. We conduct experiments on ShapeNet In-Context Datasets using PIC-G. We report the results of denoising and part segmentation.
  Models Joint Sampling
Denoising
@@
Part Segmentation
@@
PIC-G-Cat 29.3 17.03
5.3 78.95
PIC-G-Sep 36.3 23.72
7.6 74.95
TABLE IV: Influence of prompt selecting strategies. We explore the influence of prompt quality on the model’s performance. We conduct comparative experiments using PIC-G and PIC-S on the ShapeNet In-Context Datasets and Human & Object Segmentation Datasets, respectively. Note that AKB-48 is not included in the training set. Gray indicates our default settings in Tab. I and Tab. II.
(a) Ablations of PIC-G on ShapeNet In-Context Datasets.
  Models
Selection
method
Reconstruction
[…]
Denoising
[…]
Registration
[…]
PIC-G-Cat Random 4.3 5.3 14.1
@@
(b) Ablations of PIC-S on Human & Object Segmentation Datasets.
  Models
Selection
method
@@
TABLE V: Ablation study on mask ratio. We conduct ablation experiments on the separated version of PIC-G (PIC-G-Sep) on the ShapeNet In-Context Datasets. For reconstruction, denoising, and registration, we report the CD loss (x1000). For part segmentation, we report the mIoU metric.
Mask
Ratio
Reconstruction
[…]
Denoising
[…]
Registration
[…]
Part Seg.
@@
5.3 Results on Human & Object Segmentation
Comparison to Other Methods. We compare our PIC-G and PIC-S with other methods, all of which have well-designed model architectures and segmentation task heads. As shown in Tab. II, the part segmentation performance of PIC-G under multi-dataset joint training is suboptimal. At the same time, PIC-S, the extended version of PIC-G, surpasses all other models and achieves SOTA results on Human & Object Segmentation Datasets. This result indicates that our extended version, PIC-S, is more adept at integrating multiple datasets in segmentation tasks.
As shown in the last column of Tab. II, our PIC-S not only achieves SOTA results on the in-domain datasets (ShapeNetPart, Human3D, BEHAVE), but also exhibits strong generalization capabilities on out-of-domain datasets (AKB-48). PIC-S surpasses PIC-G and other models by a large margin, showing superior generalization performance.
Comparison to PIC-G. Although PIC-G performs well on multitask benchmarks, it falls slightly short on segmentation tasks. Firstly, due to the increasing number of cumbersome label points with the growing categories of point cloud parts, it struggles to handle multi-dataset joint training effectively. Secondly, under the circumstance of fixed label points, it fails to learn important mapping relationships from prompts during testing on out-of-domain datasets. In contrast to PIC-G, our proposed extension, PIC-S, handles multi-dataset joint training well and demonstrates robust performance. Additionally, with the training strategy of randomly assigning labels, our PIC-S is not confined to pre-defined fixed label points but learns the mapping relationship within the real-time prompts. This enables our PIC-S to generalize to out-of-domain datasets and achieve excellent performance.
We also visualize the predictions of PIC-S-Sep and PIC-G-Sep on the Human & Object Segmentation Datasets, along with their corresponding target point clouds. To facilitate better observation, we assign colors to the model’s output results to distinguish different parts. As shown in Fig. 8, PIC-S segments the point clouds more accurately than PIC-G. Since the label-point assignment methods are different between PIC-S and PIC-G, for convenience, we only ensure that the same parts between the output point cloud and the target point cloud are visualized using the same color.
More Generalization Capability. We further explore the generalization capability of PIC-S. Compared to the fixed label points of PIC-G, PIC-S appears to be more flexible in segmentation tasks. We customize prompts to specify part segmentation for query point clouds. As shown in Fig. 9, we combine the parts of target point clouds in the prompt into unseen part division categories and utilize this customized prompt to guide the model in performing the same part segmentation. Our PIC-S can accurately generate unique part segmentation results via customized prompts.
5.4 Ablation Study
TABLE VI: Ablation study on PIC-G. We conduct ablation experiments on ShapeNet In-Context Datasets. Gray: default setting.
@@ 5.4 Ablation Study
5.4.1 Ablations on PIC-G
Effectiveness of Joint Sampling. We investigate whether the JS module has a valid effect on four tasks. As shown in Tab. III, both PIC-G-Sep and PIC-G-Cat face a sharp decline in performance on the four tasks when the JS module is absent, and they even fail to achieve the most basic goal of reconstructing masked tokens. These findings validate our intuition that maintaining the consistency of the input and target token sequence positions is an indispensable design. That is to say, the JS module is a simple yet effective way to compensate for the missing positional information.
Influence of Prompt Selecting Strategies. We explore the impact of the prompt selection on the model’s prediction results. For the random selection method, during testing, we randomly select a pair of input-target point clouds from the training set that performs the same task as the query point cloud, serving as a task prompt. For the class-aware selection method, based on the previous one, we further select point clouds belonging to the same category as the query, such as airplanes, tables, chairs, etc. Additionally, we further delve into selecting two alternative prompts. To select pairs of examples that are paired with the query point cloud, we consider two factors: the Chamfer Distance (CD) [104] between the prompt and the query point cloud, and the feature similarity between them (features are extracted from pre-trained PointNet [38]), which are respectively denoted as CD-aware and Fea-aware. As depicted in Tab. IV(b) (a)(b), the CD-aware method demonstrates the best performance and even outperforms the task-specific models on the registration task. Note that we report the random method in the main results (Tab. I), which means our model has a higher ceiling. This provides us a great opportunity to improve downstream task results by selecting higher-quality prompts, which will be the direction of future work.
[…]

### v2

@@ 5.1 Experimental Details
Table 4: Comparison results on Multi-Entity Segmentation In-Context Datasets, including ShapeNetPart [shapenetpart], Human3D [human3d], and BEHAVE [behave]. Additionally, we conduct one-shot generalization testing experiments on AKB-48 [akb48], which is not included in the training set. $\dagger$ represents training from scratch, $\ddagger$ represents full fine-tuning. We report the instance mIoU metric.
@@
$\uparrow mIoU\uparrow$ $\uparrow mIoU\uparrow$ $\uparrow mIoU\uparrow$
@@
PointNet [pointnet] CVPR’17 77.45 80.12 81.05 83.91 78.63 /
DGCNN [dgcnn] TOG’19 76.16 78.23 74.57 76.28 86.18 /
PCT [pct] CVM’21 79.46 83.71 81.65 83.16 79.42 /
@@
PointNet [pointnet] CVPR’17 75.77 79.07 77.44 79.33 66.06 41.38
DGCNN [dgcnn] TOG’19 75.16 78.58 70.52 74.55 77.48 36.07
PCT [pct] CVM’21 79.65 82.00 79.46 80.75 72.41 40.17
PTv1 [pointtransformer] ICCV’21 80.12 83.24 79.64 81.11 77.46 36.47
PointMLP [pointmlp] ICLR’22 83.76 86.32 81.34 84.64 84.56 30.86
PointNeXt [pointnext] NeurIPS’22 79.87 82.92 82.06 84.52 76.97 37.39
PTv2 [pointtransformerv2] NeurIPS’22 81.74 83.18 80.84 82.15 82.63 33.67
PointM2AE [pointm2ae] NeurIPS’22 77.37 79.46 76.20 78.36 78.31 39.30
ACT [act] ICLR’23 84.43 85.98 82.58 84.69 88.33 34.69
I2P-MAE [i2pmae] CVPR’23 74.84 77.68 74.36 78.57 79.00 16.89
MAMBA3D [han2024mamba3d] MM’24 82.26 83.51 78.92 82.36 82.27 42.25
PointDiff [pointdiff] CVPR’24 80.64 82.14 80.32 83.04 80.88 37.36
PTv3 [pointtransformerv3] CVPR’24 82.37 82.45 81.04 83.13 85.63 44.42
PointLoRA [pointlora] CVPR’25 76.17 78.47 76.17 79.02 76.16 37.27
PCMamba [pcm] AAAI’25 81.39 83.89 81.22 83.74 83.64 38.18
@@
”-Cat” PIC NeurIPS’23 64.52 70.50 72.79 75.28 83.66 9.97
PIC++ / 81.13 85.32 77.33 79.13 86.22 73.11
PIC++ / 84.13 87.31 80.11 83.34 88.03 73.52
”-Sep” PIC NeurIPS’23 79.22 81.72 74.74 77.93 84.03 19.58
PIC++ / 80.64 83.26 80.39 82.36 87.74 68.96
PIC++ / 85.14 87.82 82.82 85.59 88.63 63.08
3) Multitask Models. For PointNet [pointnet], DGCNN [dgcnn], and PCT [pct], we utilize a pre-trained backbone that is trained on the ShapeNet [7] dataset for classification tasks. This pre-trained backbone network is equipped with multiple task-specific heads to perform multitask learning on our benchmark. PointMAE [pointmae] is a masked auto-encoder. Unlike Point-BERT, PointMAE directly rebuilds points in each local area. ACT [act], I2P-MAE [i2pmae], and ReCon [recon] are recent SoTA methods that involve other modalities like image and text knowledge in the pre-training stage and enhance the performance on different tasks after fine-tuning the models. Similar to multitask models, we combine a pre-trained encoder with different task heads for training on the four tasks.
@@ 5.2 Results on Point Cloud Multitasking
Figure 8: Visualization of comparison results between PIC++(extended version) and PIC (vanilla version). PIC++ generates more accurate segmentation results than PIC, as can be seen from the spots with red arrows ($\searrow$) pointed to them.
We report extensive experimental results of various models on the dataset we proposed in Tab. 2. From this, we find that our PICs exhibits impressive results and are capable of adapting to different tasks after only one training, achieving SoTA results in all four tasks. Besides, we visualize the in-context 3D inference results of the ”-Sep” variant of PIC in Fig. 5, where our model can generate the corresponding predictions given the provided prompts among all tasks, including reconstruction, denoising, registration, and part segmentation.
Comparison to Task-Specific Models. Task-specific models, including PointNet [pointnet], DGCNN [dgcnn], PCT [pct], and ACT [act] outperform PIC in most indicators on all tasks. However, in the part segmentation task, PIC++ achieves better results than PointNet, DGCNN, PCT, and ACT, though they are trained for segmentation. It needs to be clarified that the performance of our model is directly related to the choice of prompts. When the quality of the prompt is better, PIC can achieve much better results, and this will be demonstrated in the following section.
@@
Figure 9: Generalization of PIC++. We use customized prompts to guide the model to perform specified part segmentation. The red boxes indicate the output of the ”-Sep” variant of PIC++.
@@
Comparison to Multitask Models. As shown in Tab. 2, our PIC and PIC++ outperform multitask models in multitasking learning, especially in part segmentation, while the multitask models suffer from confusion from multiple types of tasks during training. We also compare visualization results between our PIC models and multitask models on three tasks, including reconstruction, denoising, and registration. As shown in Fig. 6, our PIC-Sep and PIC-Cat outputs are more satisfactory than other multitask models.
@@
Generalization. Usually, in-context learning has a certain degree of generalization ability, thus allowing the model to adapt to different tasks quickly. This also applies to our models. First, we test the proposed methods on out-of-distribution point clouds by conducting a registration evaluation on the ModelNet40 [modelnet40] dataset. As shown in Fig. 7 (a), our models, PIC-Sep and PIC-Cat, both exhibit superior performance on open-class tasks compared to supervised learning models specifically trained for the registration task. We compare it with models trained on the single task, such as PointMAE [pointmae], and PCT [pct]. These models suffer obvious performance drops when transferred to the new dataset. The greater the difficulty, the greater the decline in performance.
@@
Training on Large-Scale Objaverse [objaverse]. To assess the impact of data scale on the model, we incorporate a large-scale dataset, Objaverse, into the training set. This dataset comprises 800K point clouds, with each sample including only XYZ coordinates. During training, we construct input-target pairs for Objaverse point clouds in the forms of point cloud reconstruction, denoising, and registration, and perform joint training with the original training data. As shown in Tab. 3(b) (a), the performance of PIC improves significantly after incorporating Objaverse. Additionally, we test the generalization of the above model on ModelNet40. As shown in Tab. 3(b) (b), the ability of PIC to generalize to unseen datasets improves considerably after adding the large-scale data. Overall, these results indicate that increasing the data scale helps to enhance the model’s in-context learning ability.
Table 5: Effectiveness of Joint Sampling Module. We conduct experiments on ShapeNet In-Context Datasets using PIC. We report the results of denoising and part segmentation.
@@
”-Cat” 29.3 17.03
5.3 78.95
”-Sep” 36.3 23.72
7.6 74.95
5.3 Results on Multi-Entity Segmentation
Comparison to Other Methods. We compare our PIC and PIC++ with other methods, all of which have well-designed model architectures and segmentation task heads. As shown in Tab. 4, the part segmentation performance of PIC under multi-dataset joint training is suboptimal. To ensure a fair comparison, we additionally conduct experiments on PIC++ trained from scratch (denoted by $\dagger$). As illustrated in the table, both versions of PIC++ achieve competitive results. Specifically, the scratch-trained PIC++ consistently outperforms the baseline PIC, while the fine-tuned PIC++ (denoted by $\ddagger$) surpasses all other models and achieves SoTA results on Multi-Entity In-Context Datasets. Whether in in-domain or out-of-domain datasets, this result indicates that our extended version, PIC++, is more adept at integrating multiple datasets in segmentation tasks.
As shown in the last column of Tab.4, our PIC++ not only achieves SoTA results on the in-domain datasets(ShapeNetPart, Human3D, BEHAVE), but also exhibits strong generalization capabilities on out-of-domain datasets (AKB-48). Specifically, PIC++ trained from scratch achieves impressive generalization performance (73.11% for -Cat” and 68.96% for-Sep”), which is comparable to the fine-tuned version’s 73.52% (-Cat”) and even superior to its 63.08% (-Sep”). Overall, PIC++ surpasses PIC and other models by a large margin, showing superior generalization performance.
Table 6: Influence of prompt selecting strategies. We explore the influence of prompt quality on the model’s performance. We conduct comparative experiments using PIC and PIC++ on the ShapeNet In-Context Datasets and Multi-Entity In-Context Datasets, respectively. Note that AKB-48 is not included in the training set. Gray indicates our default settings in Tab. 2 and Tab. 4.
(a) Experimental results of PIC on ShapeNet In-Context Datasets.
@@ 5.3 Results on Multi-Entity Segmentation
”-Cat” Random 4.3 5.3 14.1
@@
(b) Experimental results of PIC++ on Multi-Entity In-Context Datasets.
Variants Selection method
@@
Table 7: Ablation study on mask ratio. We conduct ablation experiments on the separated version of PIC (”-Sep”) on the ShapeNet In-Context Datasets. For reconstruction, denoising, and registration, we report the CD loss (x1000). For part segmentation, we report the mIoU metric.
@@
Table 8: Ablation study on PIC. We conduct ablation experiments on ShapeNet In-Context Datasets. Gray: default setting.
@@
Comparison to PIC. Although PIC performs well on multitasking benchmarks, it falls slightly short on segmentation tasks. Firstly, due to the increasing number of cumbersome label points with the growing categories of point cloud parts, it struggles to handle multi-dataset joint training effectively. Secondly, under the circumstance of fixed label points, it fails to learn important mapping relationships from prompts during testing on out-of-domain datasets. In contrast to PIC, our proposed extension, PIC++, handles multi-dataset joint training well and demonstrates robust performance. The fact that the scratch-trained PIC++ significantly outperforms the baseline PIC on all metrics (e.g., improving AKB-48 mIoU from 19.58% to 68.96% in the “-Sep” setting) strongly validates the effectiveness of our In-Context Labeling (ICL) and In-Context Enhancing (ICE) strategies. Additionally, with the training strategy of randomly assigning labels, PIC++ is not confined to pre-defined fixed label points but learns the mapping relationship within the real-time prompts. This enables PIC++ to generalize to out-of-domain datasets and achieve excellent performance.
We also visualize the predictions of the ”-Sep” model of PIC++ and PIC on the Multi-Entity In-Context Datasets, along with their corresponding target point clouds. To facilitate better observation, we assign colors to the model’s output results to distinguish different parts. As shown in Fig. 8, PIC++ segments the point clouds more accurately than PIC. Since the label-point assignment methods are different between PIC++ and PIC, for convenience, we only ensure that the same parts between the output point cloud and the target point cloud are visualized using the same color.
More Generalization Capability. We further explore the generalization capability of PIC++. Compared to the fixed label points of PIC, PIC++ appears to be more flexible in segmentation tasks. We customize prompts to specify part segmentation for query point clouds. As shown in Fig. 9, we combine the parts of target point clouds in the prompt into unseen part division categories and utilize this customized prompt to guide the model in performing the same part segmentation. Our PIC++ can accurately generate unique part segmentation results via customized prompts.
Figure 10: Robustness on different resolutions and scales.
Robustness of PIC++. We further explored the robustness of PIC++ under various perturbations. Specifically, we subjected the query point clouds to random point dropping (simulating varying resolutions) and random scale transformations, while using the same pair of prompt point clouds to guide the part segmentation task. As illustrated in Fig. 10, our model maintains excellent inference performance despite these perturbations, demonstrating the strong robustness of PIC++. Furthermore, we applied random rotations to the query point clouds while keeping the prompt conditions unchanged. As shown in Fig. 11, PIC++ retains consistently high segmentation performance. These results indicate that our PIC++ does not rely on spatial alignment between the prompt and query information, but rather effectively leverages the semantic alignment of parts.
5.4 Ablation Study
Figure 11: Robustness on rotation.